OpenStudy (anonymous):

Find all values of x between 0 and 4pi Cos(x)= -0.6591

OpenStudy (anonymous):

Could someone please explain it to me. Thank you

OpenStudy (anonymous):

use ur calculator and find Cos^-1 -0.6591 cause cos <0 so im sure that X will be at Quadrant II or III

OpenStudy (anonymous):

I did that but there are 3 other values I need to find within the 0 and 4 pi restriction

OpenStudy (anonymous):

How can I find the other three?

OpenStudy (anonymous):

What is the rule I need to follow? Cheers

ganeshie8 (ganeshie8):

\(\large \mathbb{\cos (x) = \cos(x + 2\pi ) }\)

ganeshie8 (ganeshie8):

that gives u 2nd solution

OpenStudy (anonymous):

I don't get it

ganeshie8 (ganeshie8):

wat did u get for first value ?

OpenStudy (anonymous):

Where did you come up with cosx = cos(x+2pi)

ganeshie8 (ganeshie8):

from many places. for now u can think of it like this : period of cosx = 2pi, so it repeats every 2pi.

OpenStudy (anonymous):

I got 2.290

ganeshie8 (ganeshie8):

yes, use the given property now. second solution = \(2.29 + 2\pi\)

ganeshie8 (ganeshie8):

if x is a solution, then x+2pi is also a solution. cuz cos(x) repeats every 2pi.

ganeshie8 (ganeshie8):

if that makes any sense..

OpenStudy (anonymous):

I see what about the other two values?

OpenStudy (anonymous):

It does

ganeshie8 (ganeshie8):

good :) for other two values, use another property : \(\large \cos(x) = \cos(-x)\)

ganeshie8 (ganeshie8):

let me show it in a picture, so it makes sense why this is true.

ganeshie8 (ganeshie8):

u familiar wid 'unit circle' definition of trig functions ?

OpenStudy (anonymous):

Yep

ganeshie8 (ganeshie8):

then its easy for u,

ganeshie8 (ganeshie8):